OpenStudy (anonymous):

help simplifying (5p^3x^-2y^4z^-6 \ 15p^0x^-9y^-6z^-8) ^3 please :)

OpenStudy (amistre64):

\[\frac{5p^3x^-2y^4z^-6}{15p^0x^-9y^-6z^-8) ^3}\] this seems typoed?

OpenStudy (amistre64):

\[(\frac{5p^3x^{-2}y^4z^{-6}}{15p^0x^{-9}y^{-6}z^{-8}})^3\] this seems typoed?

OpenStudy (amistre64):

5/15 = 1/3 p^(3-0) = p^3 x^(-3--9) = x^6 z^(-6--8) = z^2

OpenStudy (amistre64):

\[(\frac{p^3 x^6 z^2}{3})^3 = \frac{p^9 x^{18} z^6}{9}\]

OpenStudy (anonymous):

it wasnt typoed lol that was the actual problem.